# High-temperature domain wall current in Mg-doped lithium niobate single crystals up to 400C

### v1

@@ Abstract
Conductive ferroelectric domain walls (DWs) represent a promising topical system for the development of nanoelectronic components and devices. DWs may show very different properties as compared to their bulk counterparts. Of central interest here is the domain wall current (DWC) of charged DWs in 5molMg-doped lithium niobate single crystals; in contrast to former works, we extend the DWC study here to temperatures as high as 400C. Both the temporal stability and the thermal activation energies of 100 - 160 meV are readily deduced from current-voltage sweeps as recorded over multiple heating cycles. Our experimental work is backed up by atomistic modelling of the DWC. The latter suggests that a large band bending renders head-to-head (H2H) and tail-to-tail (T2T) DWs semimetallic. These detailed investigations underline the potential to extending DWC-based nanoelectronic applications even into the so-far unexplored high-temperature regime.
@@
One major key factor that contributes to the versatility of LN-based devices is the ability to elegantly engineer ferroelectric domains and domain walls (DWs) into these crystals Meier and Selbach (2021); McCluskey et al. (2022); Kirbus et al. (2019); Geng et al. (2021); Waser and Rüdiger (2004). Domain engineering in LN involves the controlled creation and manipulation of DWs Shur, Akhmatkhanov, and Baturin (2015). Recently, considerable advancements have been reported for fabricating and tuning the DW properties in LN crystals, specifically also to reproducibly increase the DW conductivity (DWC) by several orders of magnitude Werner et al. (2017); Godau et al. (2017); Zhang et al. (2022); Kirbus et al. (2019), which is of clue interest to the work here. Controlling the DWC has a broad impact for various devices, such as resistive switches, diode structures, or neuromorphic computing Zhang et al. (2021); Sharma and Seidel (2023); Chai et al. (2021); Suna et al. (2022). For sensing devices, the ability to manipulate the DWC in LN offers promising prospects, provided they meet the challenging environmental conditions of i.e. high temperatures, as requested for instance in aerospace and automotive applications Turner, Fuierer, and Newnham (1994). In fact, despite some minor hints towards a thermally-activated DWC process up to 70C, studies on this topic have only rarely been reported Werner et al. (2017); Zahn et al. (2024) and the high temperature behavior and stability of domains and DWs remains unclear to date. While domains in LN endure annealing up to the Curie temperature $T_{c}$ at 1100C, the stability and enhanced conductivity of DWs is uncertain due to their non-equilibrium, inclined structure Godau et al. (2017).
@@
To clarify these issues, we (i) investigate the DWC in LN at elevated temperatures up to 400C. This allows to proving the existence of thermally-activated processes beyond room temperature, and hence to discriminating between the two mechanisms proposed above. Furthermore, (ii) we present DFT-calculated band structures of fully charged DWs in LN, in order to support whether or not the 2DEG theory accounts also for the DWC as observed here in LN-DWs.
@@
In our experiments, z-cut 5molMg-doped LN wafers from Yamaju Ceramics Co., Ltd., are used for domain engineering. Every sample piece measures 200 $\mu$m in thickness, and has a lateral size of 5 mm $\times$ 6 mm. Single, hexagonally-shaped domains of a 110 - 300 $\mu$m diameter are fabricated by means of our UV-assisted electric field poling technique Haußmann et al. (2009); Godau et al. (2017). In total, we investigated three different samples in this study, tagged as LN1, LN2, and LN3. Note that sample LN1 carries one such poled domain only, while two broadly separated domains $a$ and $b$ were poled into samples LN2 and LN3; LN2-a and LN2-b hence refer to the domain $a$ and $b$ in sample LN2, respectively. This variation aims at exploring potential relationships between domain size, current flow, and thermal stability. Domain formation is monitored $in-situ$ via polarized light microscopy. Individual domains are then electrically contacted by depositing a 10-nm-thick Cr-electrode via magnetron sputtering onto both z-facets, and wired up to our electrical measuring units through conductive silver-paste contacts. As-grown DWs typically have a very low conductivity at room temperature (RT) that then is enhanced by several orders of magnitude following the protocol of Godau et al.Godau et al. (2017); Kirbus et al. (2019).
DWC investigations in the temperature range of 25-400C are then carried out by mounting the sample onto a ceramic heating element. Current-voltage (I-V) sweeps between ±10 V are recorded using a Keithley 6517B electrometer connected to a $\mu$-impedance setup, applying voltage increments of 1.48 V and a waiting time per step of 10 seconds. Using the $\mu$-impedance setup bears two advantages, firstly in allocating individual DWs and surface spots with high precision through x-y positioning the top-contact platinum tip, while secondly, facilitating data acquisition at various temperatures considering a heating rate of 1 K/min, only.
@@
Figure 1: I-V-characteristic across the $\pm$10 V voltage bias regime of the domain wall current (DWC) in sample LN3-a. (a) DWC before (black) and after (red) conductivity enhancement at room temperature; and (b) DWC recorded at elevated temperatures up to 400C. Note the ohmic behavior of all I-V-curves, being independent on temperature or domain size. The inset in (a) depicts the DW contour of sample LN3-a, as recorded by polarized light microscopy.
@@
Fig. 1 displays the typical DWC I-V characteristics for sample LN3-a at different stages and temperatures. Directly after domain poling illustrated also by the polarized microscopy image (see inset in Fig. 1a) the DWC at RT is negligibly small across the full ±10 V range (black curve in Fig. 1a), while showing a drastic increase of up to six orders of magnitude after enhancement (red curve in Fig. 1a). Note that all DWs studied in this work exhibit an ohmic behavior after enhancement, while in previous works often rectifying characteristics have been reported as well Godau et al. (2017); Kirbus et al. (2019); Zahn et al. (2024). We find absolute room temperature DWCs ranging from $\sim$10-7 to $\sim$10-6A at a $+$9 V bias voltage for all DWC measurements in all three samples, with differences being indicative for their effective and averaged DW inclination angle with respect to the sample c-axes Godau et al. (2017). Notably, the DWCs studied here show no dependence on either the DW area or size.
Fig. 1b presents the DWC I-V-curves recorded from DW LN3-a at elevated temperatures, i.e. at 100 C (green), 350 C (black), 375 C (red), and 400 C (blue). Notably all I-V-curves behave ohmic-like, while the DWC at 400 C is increased at least 600 $\times$ as compared to the RT DWC.
@@
Fig. 2 displays the DWC recorded at a $+$7 V bias voltage for samples LN1-a and LN3-a in the Arrhenius-type plot, together with the DWC data from a bulk LN reference sample containing no DW at all (plotted in orange). The DWC increases as a function of rising temperatures for all samples, including the bulk reference. This trend is observed consistently either for increasing or decreasing temperature sweeps. Of particular interest is the comparison between investigated DWCs and bulk LN currents, as depicted in Fig. 2. The bulk LN current follows from conductivity data obtained through impedance spectroscopy Yakhnevych et al. (2024); Kofahl et al. (2024); here, currents were calculated based on the sample thickness and average electrode size. Note the huge increase in the current magnitude between the bulk reference and samples LN1-a and LN3-a, respectively, displaying a six-order-of-magnitude difference at 250 C. Therefore, we can exclude any bulk related contribution to the measured DWCs.
@@
Figure 2: Temperature-dependent DWC plotted for DW-sample LN1-a (purple) and LN3-a (blue) at a +7 V bias voltage. For reference, the bulk LN current (in orange) is displayed in the same Arrhenius plot.
Moreover, we notice in Fig. 2 temperature ranges (marked with solid lines and annotated in Roman numerals: I, II, III) that show a linear increase in DWC, however, with different slopes, indicating thermally activated transport processes with different activation energies. We see from Fig. 2 that different DWC samples (here LN1-a and LN3-a) scatter by one order of magnitude in the absolutely measured DWC (see the DWC data for sample LN1-a and LN3-a at RT in Fig. 2). This trend persists up to approximately 200C, hence branding the region I in Fig. 2. Above that threshold of 200C, DWC data scatter less, (region II) and the differences gradually diminish, converging practically to the same DWC value for all analyzed DWs (in region III). These findings suggest that the DW morphology plays no major role anymore above the 200C threshold.
@@
Fig. 3 presents the activation energy $E_{A}$ for all investigated DWCs as extracted by fitting $I\sim\exp(E_{A}/k_{B}T)$ to the linear segments I, II, and III in Fig. 2. The temperature intervals used for data fitting are the same ones as the ones delineated in Fig. 2. In addition, Fig. 3 contains also the as-deduced $E_{A}$ for DW LN2-a, depicted for both its initial DWC state (solid black line) and the subsequent heating cycles 1 (green solid line) and 2 (green dashed line). Remarkably, our findings suggest that the studied domains maintain stable during repeated heating cycles. This implies a promising potential for utilizing DWC-based devices at elevated temperatures.
@@
Figure 3: Activation energies $E_{A}$ of the DWC for the distinct temperature ranges as marked in Fig. 2. Note the reproducible $E_{A}$ values of DW LN2-a even after several heating cycles.
@@
Our observations in Fig. 3 reveal two distinct activation energy levels, measuring approximately $E_{A}$ = 100 meV and 160 meV. The data analysis delineates the temperature ranging into two overlapping regimes: room to medium temperatures (25 - 280C) and temperatures above about 150C. Werner et al. Werner et al. (2017) and Zahn et al. Zahn et al. (2024) reported $E_{A}$ values of 100 meV and 200 meV for the temperature range from -193C to 70C. They also suggested that the dominant charge carrier type for DWC in these DWs is of electronic nature, backed up by both recent Hall-transport measurements Beccard et al. (2023) and strain vs. conductivity investigations Singh et al. (2022), specifically involving thermally activated electron-polaron hopping. We assume that the activation energy values obtained here, ranging from 90 meV to 160 meV, also indicate involvement of electron-polaron hopping processes; in fact, such $E_{A}$ values align well with the reported values for electron-polaron jumps in bulk lithium niobate Reichenbach et al. (2018); Koppitz, Schirmer, and Kuznetsov (1987); Schirmer et al. (2009).
@@
In order to investigate the mechanisms behind the DWC, fully-charged head-to-head (H2H) and tail-to-tail (T2T) DWs of stoichiometric LN are modeled within DFT in the independent particle approximation (IPA). We thereby employ the VASP software package Kresse and Furthmüller (1996a, b), PAW potentials implementing the PBE formulation of the xc-functional Perdew, Burke, and Ernzerhof (1996); Perdew et al. (2008); Blöchl (1994) and a cutoff energy for the plane wave basis at 500 eV. A H2H and a T2T DW are modelled within a 1$\times$1$\times$12 repetition of the hexagonal unit cell and a 4$\times$4$\times$1 k-point mesh. Thus, the two DWs, one H2H and one T2T, are separated by about 8 nm. Structural optimization reveals that the atomic positions and the local polarization reach the value of bulk LN at a distance of about 1-2 nm away from the DW center, which is in agreement with typical DW widths as deduced from TEM investigations Gonnissen et al. (2016). The local density of states of the super-cell for modelling the DW is shown in Fig. 4. The direction of polarization and the sign of the polarization charges are indicated as well. Interestingly, the spontaneous polarization causes a strong band bending for both types of fully-charged DWs, so that a non-vanishing portion of electronic states belonging to the conduction (valence) band, crosses the Fermi energy in proximity to the H2H (T2T) DW Verhoff et al. . The domain walls thus become semi-metallic, similar to suggestions in other materials Nataf et al. (2020); Bednyakov et al. (2018). Moreover, the DOS at the H2H and T2T DW is very different; H2H DWs feature very low DOS values at the Fermi-energy and higher DOS values about 150 meV below it. Unfortunately, due to the limited precision of the DFT approach, it cannot be conclusively stated whether this gap in the H2H DWs DOS is related to the measured activation energy or not.
@@
  • Firstly, improved methods beyond the IPA approach used here for calculating the electronic structure, are required to quantitatively estimate the overlap between conduction and valence bands.
  • Second, non-charged or only slighted-inclined DWs, such as the DW measured in the present experiments, feature a less pronounced band bending and hence might not lead to the formation of a 2DEG. Whether the contribution of H2H and T2T walls within an oblique DW suffices to render the whole DW metallic is questionable.
@@
In conclusion, our results consistently demonstrate that as temperature rises, there is a corresponding increase in the domain wall current (DWC). DWs, along with their conductivity, remain stable even upon multiple heating cycles up to about 400C, marking the highest temperatures documented in existing literature. This indicates the potential for developing DW-based electronics for high temperature-applications. At temperatures up to about 400C, the thermally-activated nature of the DWC is confirmed, indicating that a hopping process (or similar mechanism) is also present for this temperature range. Here, no signs of saturation or an extrinsic conductivity region are found for the slightly-inclined DWs. Furthermore, DWs display varied activation energies in two overlapping temperature regimes, hinting towards the involvement of more than one type of defect/polaron to conduction, consistently observed within several DWs.

### v2

@@ Abstract
Conductive ferroelectric domain walls (DWs) represent a promising topical system for the development of nanoelectronic components and devices. DWs show very different properties as compared to their bulk counterparts. Of central interest here is the domain wall current (DWC) of charged DWs in 5molMg-doped lithium niobate single crystals; in contrast to former works, we extend the DWC study here to temperatures as high as 400C. Both the temporal stability and the thermal activation energies of 90 - 160 meV are readily deduced from current-voltage sweeps as recorded over multiple heating cycles. Our experimental work is backed up by atomistic modelling of the DWC. The latter suggests that a large band bending renders head-to-head and tail-to-tail DWs semimetallic. These detailed investigations underline the potential to extend DWC-based nanoelectronic applications even into the so-far unexplored high-temperature regime.
@@
One major key factor that contributes to the versatility of LN-based devices is the ability to elegantly engineer ferroelectric domains and domain walls (DWs) into these crystals Meier and Selbach (2021); McCluskey et al. (2022); Kirbus et al. (2019); Geng et al. (2021); Waser and Rüdiger (2004). Domain engineering in LN involves the controlled creation and manipulation of DWs Shur et al. (2015). Recently, considerable advancements have been reported for fabricating and tuning the DW properties in LN crystals, specifically also to reproducibly increase the DW conductivity (DWC) by several orders of magnitude Werner et al. (2017); Godau et al. (2017); Zhang et al. (2022); Kirbus et al. (2019) which is of clue interest to the work here. Controlling the DWC has a broad impact for various devices, such as resistive switches, diode structures, or neuromorphic computing Zhang et al. (2021); Sharma and Seidel (2023); Chai et al. (2021); Suna et al. (2022). For sensing devices, the ability to manipulate the DWC in LN offers promising prospects, provided they meet the challenging environmental conditions of i.e. high temperatures, as requested for instance in aerospace and automotive applications Turner et al. (1994). In fact, despite some minor hints towards a thermally-activated DWC process up to 70C, studies on this topic have only rarely been reported Werner et al. (2017); Zahn et al. (2024) and the high-temperature behavior and stability of domains and DWs remain unclear to date. While domains in LN endure annealing up to the Curie temperature ($T_{c}$) at about 1150C, the stability and enhanced conductivity of DWs is uncertain due to their non-equilibrium, inclined structure Godau et al. (2017).
@@
To clarify these issues, we (i) investigate the DWC in LN at elevated temperatures up to 400C. This allows us to prove the existence of thermally-activated processes beyond room temperature, and hence to discriminate between the two mechanisms proposed above. Furthermore, (ii) we present density functional theory (DFT) calculation of band structures of fully charged DWs in LN, in order to support whether or not the 2DEG theory accounts also for the DWC as observed here in DWs of Mg-doped LN.
@@
In our experiments, z-cut 5molMg-doped LN wafers from Yamaju Ceramics Co., Ltd., are used for domain engineering. Every sample piece measures 200 $\mu$m in thickness, and has a lateral size of 5 mm $\times$ 6 mm. Single, hexagonally-shaped domains of a 110 - 300 $\mu$m diameter are fabricated by means of our UV-assisted electric field poling technique Haußmann et al. (2009); Godau et al. (2017). In total, we investigated three different samples in this study, tagged as LN1, LN2, and LN3. Note that samples LN1 and LN3 carry one such poled domain denoted by LN1-a and LN3-a, respectively, while two broadly separated domains were poled into sample LN2, as denoted LN2-a and LN2-b. This variation aims at exploring potential relationships between domain size, current flow, and thermal stability. Domain formation is monitored $in-situ$ via polarized light microscopy. Individual domains are then electrically contacted by depositing a 10-nm-thick Cr-electrode via magnetron sputtering onto both z-facets, and wired up to our electrical measuring units through conductive silver-paste contacts. As-grown DWs typically have a very low conductivity at room temperature (RT) that then is enhanced by several orders of magnitude following the protocol of Godau et al. Godau et al. (2017); Kirbus et al. (2019).
DWC is investigated in the temperature range of 25-400C using a high-temperature setup. Here, the samples are placed on a ceramic plate with a planar platinum contact. Both are installed atop a heating module. Individual domains are contacted by a platinum tip that is precisely positioned using a 3D micromanipulator. A thermocouple is mounted at a distance of approximately 1-2 mm from the sample to control the temperature with a heating rate of 1 K/min. Current-voltage ($I$-$V$) sweeps between ±10 V are recorded using a Keithley 6517B electrometer connected to the platinum tip and the planar contact while applying voltage increments of 1.5 V. Details of the setup can be found in the supplementary material (SM, sec. S2).
@@
Figure 1: $I$-$V$-characteristic across the $\pm$10 V voltage bias regime of the domain wall current in domain LN3-a. The inset in depicts the DW contour of domain LN3-a, as recorded by polarized light microscopy.
@@
Fig. S1 displays the typical DWC $I$-$V$-characteristics for domain LN3-a at different temperatures. Directly after domain poling illustrated also by the polarized microscopy image (see inset in Fig. S1) the DWC at RT is negligibly small across the full ±10 V range, approximately 10-11A at +7 V. A drastic increase of up to six orders of magnitude is evident after enhancement. Note that all DWs studied in this work exhibit an ohmic behavior after enhancement, while in previous works often rectifying characteristics have been reported as well Godau et al. (2017); Kirbus et al. (2019); Zahn et al. (2024). We find absolute room temperature DWCs ranging from $\sim$10-7 to $\sim$10-6A at a $+$7 V bias voltage for all DWC measurements in all domains, with differences being indicative for their effective and averaged DW inclination angle with respect to the sample c-axes Godau et al. (2017). Notably, the DWCs studied here show no dependence on the domaim size. Remarkably, the DWC in domain LN3-a at 400 C is increased at least 800 times as compared to the RT DWC at a +7 V bias voltage. The $I$-$V$-characteristics across the ±10 V voltage bias range for the DWC in all examined domains at RT, along with polarized microscopy images illustrating the domains directly after domain poling, are available in the SM (sec. S1). Additionally, the SM includes the $I$-$V$-curve across various temperatures for all studied domains (sec. S3).
@@
Fig. S2 displays the DWC recorded at a $+$7 V bias voltage for domains LN1-a and LN3-a in the Arrhenius-type plot, together with the DWC data from a bulk LN reference sample containing no DW (plotted in orange). The DWC increases as a function of rising temperatures for all domains, including the bulk reference (see Fig. S2, Fig. S7 and Fig. S8 in SM sec. S4). This trend is observed consistently either for increasing or decreasing temperature sweeps. Of particular interest is the comparison between investigated DWCs and bulk LN currents, as depicted in Fig. S2. The bulk LN current follows from conductivity data obtained through impedance spectroscopy Yakhnevych et al. (2024); Kofahl et al. (2024); here, the current was calculated based on the sample thickness and average electrode size. Note the huge increase in the current magnitude between the bulk reference and domains LN1-a and LN3-a, respectively, displaying a six-order-of-magnitude difference at 250 C. Therefore, we can exclude any noticable bulk related contribution to the measured DWCs. Moreover, we notice in Fig. S2 different temperature regimes (marked with solid lines and annotated in Roman numerals: I, II, III) that show a linear increase in DWC, however, with different slopes, indicating thermally-activated transport processes with different activation energies. We see from Fig. S2 that different DWCs (here LN1-a and LN3-a) scatter by one order of magnitude in the absolutely measured DWC (see the DWC data for domain LN1-a and LN3-a at RT in Fig. S2). This trend persists up to approximately 200C, hence branding the region I in Fig. S2. Above that threshold of 200C, DWC data scatter less, (region II) and the differences gradually diminish, converging practically to the same DWC value for all analyzed DWs (in region III). These findings suggest that the DW morphology plays no major role anymore above the 200C threshold.
@@
Figure 2: Temperature-dependent DWC plotted for domain LN1-a (purple) and LN3-a (blue) at a +7 V bias voltage. For reference, the bulk LN current (in orange) is displayed in the same Arrhenius plot.
@@
Fig. S3 presents the activation energy ($E_{A}$) for all investigated DWCs as extracted by fitting $I\sim\exp(E_{A}/k_{B}T)$ to the linear segments I, II, and III in Fig. S2. Here, $k_{B}$ represents the Boltzmann constant. The temperature intervals used for data fitting are the same ones as the ones delineated in Fig. S2. In addition, Fig. S3 contains also the as-deduced $E_{A}$ for DW LN2-a, depicted for both its initial DWC state (green solid line) and the subsequent heating cycle (green dashed line). Remarkably, our findings suggest that the studied domains maintain stable during repeated heating cycles. This implies a promising potential for utilizing DWC-based devices at elevated temperatures.
@@
Figure 3: Activation energies $E_{A}$ of the DWC for the distinct temperature ranges as marked in Fig. S2. Note the reproducible $E_{A}$ values of DW LN2-a even after several heating cycles.
@@
Our observations in Fig. S3 reveal two distinct activation energy levels, measuring approximately $E_{A}$ = 90 meV and 160 meV. The data analysis delineates the temperature ranging into two overlapping regimes: room to medium temperatures (25 - 280C) and temperatures above about 150C. Werner et al. Werner et al. (2017) and Zahn et al. Zahn et al. (2024) reported $E_{A}$ values of 100 meV and 200 meV for the temperature range from -193C to 70C. They also suggested that the dominant charge carrier type for DWC in these DWs is of electronic nature, backed up by both recent Hall-transport measurements Beccard et al. (2023) and strain vs. conductivity investigations Singh et al. (2022), specifically involving thermally activated electron-polaron hopping. We assume that the activation energy values obtained here, ranging from 90 meV to 160 meV, also indicate involvement of electron-polaron hopping processes; in fact, such $E_{A}$ values align well with the reported values for electron-polaron jumps in bulk lithium niobate Reichenbach et al. (2018); Koppitz et al. (1987); Schirmer et al. (2009).
@@
In order to investigate the mechanisms behind the DWC, DWs perpendicular to the crystallographic directions X, Y and Z of stoichiometric LN are modeled within DFT in the independent particle approximation (IPA). We thereby employ the VASP software package Kresse and Furthmüller (1996a, b), projector augmented-wave (PAW) potentials implementing the Perdew-Burke-Ernzerhof (PBE) formulation of the xc-functional Perdew et al. (1996, 2008); Blöchl (1994) and a cutoff energy for the plane wave basis of 500 eV. Further computational details can be found in the SM (sec. S5). Structural optimization reveals that the atomic positions and the local polarization reach the value of bulk LN at a distance of about 1 nm from the DW center in the case of X and Y DWs, and at a distance of 2 nm in the case of the fully-charged Z walls. This is in agreement with typical DW widths as deduced from transmission electron microscopy investigations Gonnissen et al. (2016). The presence of DWs deeply modifies the electronic structure, as shown by the calculated density of states (DOS) shown in the SM (Figure S9). However, while X and Y DWs reduce the fundamental bandgap from 3.4 eV to about 3.0 eV, charged Z DWs lead to an overlap of valence and conduction band, as demonstrated by the local density of states of the super-cell modelling the DWs shown in Fig. 4. The direction of polarization and the sign of the polarization charges for the head-to-head (H2H) and the tail-to-tail (T2T) DWs are indicated as well. Interestingly, the spontaneous polarization causes a strong band bending for both types of fully-charged DWs, so that a non-vanishing portion of electronic states belonging to the conduction (valence) band, crosses the Fermi energy in proximity to the H2H (T2T) DW Verhoff et al. . The domain walls thus become semi-metallic, similar to suggestions in other materials Nataf et al. (2020); Bednyakov et al. (2018). Moreover, the DOS at the H2H and T2T DWs is very different; H2H DWs feature very low DOS values at the Fermi-energy and higher DOS values about 150 meV below it. Unfortunately, due to the limited precision of the computational approach, it cannot be conclusively stated whether this gap in the H2H DWs DOS is related to the measured activation energy or not.
@@
  • Firstly, improved methods beyond the IPA approach as used here for calculating the electronic structure, are required to quantitatively estimate the overlap between conduction and valence bands.
  • Second, inclined DWs, such as the DW measured in the present experiments, can be considered as a sequence of non charged X and Y DWs and charged H2H and T2T walls. The first feature a less pronounced band bending and hence might not lead to the formation of a 2DEG. Whether the contribution of H2H and T2T walls within an oblique DW suffices to render the whole DW metallic is questionable. This situation will not drastically change at the temperatures discussed in this work, when vibrational contributions are expected to close the electronic bandgap by further 0.5 eV with respect to the 0 K calculations presented here Riefer et al. (2016).
@@
In conclusion, our results consistently demonstrate that as temperature rises, there is a corresponding increase in the domain wall current. DWs, along with their conductivity, remain stable even upon multiple heating cycles up to about 400C, marking the highest temperatures documented in the existing literature. This indicates the potential for developing DW-based electronics for high-temperature applications. At temperatures up to about 400C, the thermally-activated nature of the DWC is confirmed, indicating that a hopping process (or similar mechanism) is also present for this temperature range. Here, no signs of saturation or an extrinsic conductivity region are found for slightly-inclined DWs. Furthermore, DWs display varied activation energies in two overlapping temperature regimes, hinting towards the involvement of more than one type of defect/polaron to conduction, consistently observed within several DWs.